OpenStudy (amilapsn):

A punching machine has to carry out piercing operations at the rate of 12 holes per minute. The amount of work to be done in cutting 30mm diameter holes in 15mm thick plate is 8J/mm2 of the sheared area. The actual piercing operation takes 2s. The total frictional losses over a cycle are equivalent to one sixth of the work done during a single piercing operation. The machine shaft is driven by a constant power electric motor. The maximum fluctuation of the speed of the machine shaft is between 190 rev/min and 210 rev/min. (a) Find the minimum required motor power to drive the piercing machine

OpenStudy (amilapsn):

(b) Determine the required sum of the mass moments of inertia of all revolving parts on the machine shaft in order to keep the speed fluctuation within the given limits. State any assumption made.

OpenStudy (amilapsn):

@ganeshie8

OpenStudy (amilapsn):

@dan815

OpenStudy (dan815):

:)

OpenStudy (amilapsn):

I've got some problems to sort out regarding this problem @dan815

OpenStudy (dan815):

okay what seems to be the problem

OpenStudy (amilapsn):

During a piercing operation the shear area would be \(\pi r^2\) or \(s2\pi r h\) where r is the diameter of the hole while h is the depth...

OpenStudy (amilapsn):

*\(2\pi rh\)

OpenStudy (dan815):

why 2*pirh

OpenStudy (amilapsn):

Then why the depth is given?

OpenStudy (dan815):

not sure why, i dont think we need it

OpenStudy (dan815):

it gives work per surface area though in mm^2 so

OpenStudy (amilapsn):

Work done per hole = 8 x pi x r^2?

OpenStudy (amilapsn):

Then another q:

OpenStudy (dan815):

yep

OpenStudy (amilapsn):

Am I right: minimum power to drive the piercing machine=maximum power in a cycle

OpenStudy (amilapsn):